# Estimating invertible processes in Hilbert spaces, with applications to functional ARMA processes

### v2

@@ A.1 Proofs of results in Section 3
$\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\hat{\mathscr{C}}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1}\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\hat{\lambda}_{j}+\theta_{\!N})^{-1}=(\hat{\lambda}_{K}+\theta_{\!N})^{-1},$ (A.6)
$\Big{\|}\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\mathscr{C}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1}\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\lambda_{j}+\theta_{\!N})^{-1}=(\lambda_{K}+\theta_{\!N})^{-1}.$ (A.7)
@@
where the series on the right is finite by Assumption 3.2. Finally, from the definition of the H-S norm, with $\Psi_{\!L}=\Psi_{\!L}(K)$ in Assumption A.1, $\mathscr{C}^{\ddagger}=\mathscr{C}(\mathscr{C}+\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{-1},$ and the Kronecker delta $\delta_{jk},$
@@
Altogether, plugging (A.6)–(A.7) and (A.1)–(A.12) in (A.5), and using $\|\mathscr{D}\|_{\mathcal{S}}\leq\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2}$ by (A.2), $(\lambda_{K}+\theta_{\!N})^{-1}\leq\lambda^{-1}_{K}$ and $(\hat{\lambda}_{K}+\theta_{\!N})^{-1}\leq\hat{\Lambda}_{K_{b}-1},$ for each $K,L,N,$ we obtain
@@ A.3 Proofs of results in Section 5
To derive the asymptotic result for our fARMA(1,1)11(1,1)( 1 , 1 ) operators, we derive a preliminary result based on the following assumption.
@@
Suppose $(X_{k})$ is the fARMA(1,1)11(1,1)( 1 , 1 ) process in (5.2). Further, let the conditions of Theorem 3.1, Assumptions 5.1, 5.35.4, A.2, and $\gamma_{N}=\mathrm{O}(K^{-\beta})$ hold. Then,
@@
Let $(X_{k})$ be the fARMA$(p,q)$ process in (5.1). Moreover, let the conditions of Theorem 3.1, Assumptions 5.5, 5.75.8, A.3 and $\gamma_{N}=\mathrm{O}(K^{-\beta})$ hold. Then,
@@
Further, as in the proof of Lemma A.4, together with (A.20) and Assumptions 5.75.8, it holds

### v3

@@ Appendix A Appendix
Proof.
See Weidmann (1980), Theorem 7.8. ∎
The following lemma follows from elementary norm inequalities and the details are omitted.
@@
$\Big{\|}\hat{\mathscr{C}}^{\dagger}\!\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\hat{\mathscr{C}}+\hskip-0.50003pt\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{\hskip-0.40974pt-1}\coprod^{\hat{c}_{K}}_{\hat{c}_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\hat{\lambda}_{j}+\theta_{\!N})^{-1}=(\hat{\lambda}_{K}+\theta_{\!N})^{-1},$ (A.6)
$\Big{\|}\mathscr{C}^{\dagger}\!\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\Big{\|}(\mathscr{C}+\hskip-0.50003pt\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{\hskip-0.40974pt-1}\coprod^{c_{K}}_{c_{1}}\Big{\|}_{\mathcal{L}}=\sup_{j\leq K}(\lambda_{j}+\theta_{\!N})^{-1}=(\lambda_{K}+\theta_{\!N})^{-1}.$ (A.7)
@@
where the series on the right is finite by Assumption 3.2. Finally, from the definition of the H-S norm, with $\Psi_{\!L}=\Psi_{\!L}(K)$ in Assumption A.1, $\mathscr{C}^{\ddagger}=\mathscr{C}(\mathscr{C}+\hskip-0.50003pt\theta_{\!N}\hskip 0.50003pt\mathbb{I})^{\hskip-0.40974pt-1},$ and the Kronecker delta $\delta_{jk},$
@@
Altogether, plugging (A.6)–(A.7) and (A)–(A.12) in (A.5), and using $\|\mathscr{D}\|_{\mathcal{S}}\leq\sqrt{L}\operatorname{\mathds{E}}\!\|X_{0}\|^{2}$ by (A.2), $(\lambda_{K}+\theta_{\!N})^{-1}\leq\lambda^{-1}_{K}$ and $(\hat{\lambda}_{K}+\theta_{\!N})^{-1}\leq\hat{\Lambda}_{K_{b}-1},$ for each $K,L,N,$ we obtain
@@
To derive the asymptotic result for our fARMA(1,1)(1,1)( 1 , 1 ) operators, we derive a preliminary result based on the following assumption.
@@
Suppose $(X_{k})$ is the fARMA(1,1)(1,1)( 1 , 1 ) process in (5.2). Further, let the conditions of Theorem 3.1, Assumptions 5.1, 5.3, A.2, and $\gamma_{N}=\mathrm{O}(K^{-\beta})$ hold. Then,
@@
[…]
@@
Further, as in the proof of Lemma A.4, together with (A.20) and Assumption 5.6, it holds
@@
Acknowledgements
The authors are grateful to the editor, the associate editor and the two referees for their constructive comments on an earlier version of this article, which greatly improved the quality of the paper. Parts of the paper were written while Sebastian Kühnert was employed at the University of California, Davis.
Funding
Alexander Aue was partially supported by NSF DMS 2515821. Sebastian Kühnert was partially supported by TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (Project number 520388526) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation). Gregory Rice was partially supported by the Discovery Grant RGPIN 50503-11525 3100 105 from the Natural Science and Engineering Research Council of Canada.